OpenStudy (anonymous):

PLEASE HELP? SERIOUSLY? :) Solve 2x + 5y = -133x - 4y = -8 (4, 1) (-4, 1) (4, -1) (-4, -1)

OpenStudy (anonymous):

is that a set of coordinates?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

what is it ? : 2x + 5y = -133x - 4y = -8 why there are 2 equal signs

OpenStudy (anonymous):

To solve by addition, you need to determine the factors that eliminates one of the variables for each equation. E.g. 4(2x + 5y = -9) 5(3x - 4y = -2) 8x + 20y = -36 15x - 20y = -10 23x = -46 x = -2 Substitution method? Take one of the equations and solve for the variable. Then, substitute that for another equation. Simplify.

OpenStudy (anonymous):

I don't know, it's set up in the weirdest way....

OpenStudy (anonymous):

wait.. it look like that the two equations are : 2x + 5y = -9 3x - 4y = -2 ?

OpenStudy (anonymous):

Disregard the long, drawn out example i made

OpenStudy (anonymous):

ok.. its hard to understand what are you two equations from 2x + 5y = -133x - 4y = -8 it might be : 2x + 5y = -13 , 3x - 4y = -8

zepdrix (zepdrix):

Hmm that would make more sense, is that your system Chao?

OpenStudy (anonymous):

yes the OP equation dosent make sense

zepdrix (zepdrix):

ok ok ok now this is starting to make sense :)

OpenStudy (anonymous):

YES! 2x + 5y = -13 , 3x - 4y = -8

OpenStudy (anonymous):

if so .. we might multiply the first by 3 multiply the second by 2 then subtract the second from the first : 3(2x + 5y) = -13 * 3 2(3x - 4y) = -8 * 2 6x + 15y = -39 6x -8y = -16 6x + 15y -(6x-8y) = -39 - (-16) 15y + 8y = -23 23y = -23 y = -1 can you tell what is x now ?

OpenStudy (anonymous):

It's either 4, or -4

OpenStudy (anonymous):

How do I solve for x....?

OpenStudy (anonymous):

now go back to one of the original equations : for example : 2x + 5y = -13 plug in here y = -1 what you get for x ?

OpenStudy (anonymous):

If you have an x=a (a being some number) then plug a into your equation and find y, then you will have (x,y)

OpenStudy (anonymous):

What number though?

OpenStudy (anonymous):

Oh, okay!!!!! I'll use one of the numbers from the multiple choice?

OpenStudy (anonymous):

Hear back from your professor yet, @InsanelyChaotic ?

OpenStudy (anonymous):

Sorry, I know I sound like a dumb*** and get overly excited when I know something pertaining to math but bear with me

OpenStudy (anonymous):

you have to solve one of the equations for x, then substitue that value into the other.

OpenStudy (anonymous):

Yes @CliffSedge

OpenStudy (anonymous):

look at what i wrote : now go back to one of the original equations : for example : 2x + 5y = -13 plug in here y = -1 what you get for x ? if you plug y = -1 into 2x + 5y = -13 you get 2x -5 = -13 2x = -8 x= -4

OpenStudy (anonymous):

I know this much, thank you

OpenStudy (anonymous):

in fact you could, instead of solving it just plug the choices that you got up there and see which of them holds for your equations

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

I'm so confused

OpenStudy (anonymous):

I know this should be simple but it's not

OpenStudy (anonymous):

F*** math

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

what you have here are two equations that share the same value for x and y. so say you solve one equation and get x=(y with math done to it) then plug (y with math) into the other equation for x and get a value for y

OpenStudy (anonymous):

that y can be pluged back in to the OTHER equation to give you your x. once you have x and y put them in the form (x,y) and thats your answer

OpenStudy (anonymous):

I'm just getting emotional

OpenStudy (anonymous):

you got this! look up systems of linear equations in google. dont get discouraged! math makes alot more sense than emotions

OpenStudy (anonymous):

http://www.purplemath.com/modules/systlin1.htm

OpenStudy (anonymous):

I'm reading

OpenStudy (anonymous):

I'm not good at math and I'm reminded of this constantly

OpenStudy (anonymous):

I'm not good at this I'm sorry you're so nice for helping but I'm not good at any of this

OpenStudy (anonymous):

This site is also rather good: http://themathpage.com/alg/simultaneous-equations.htm

OpenStudy (anonymous):

Thank you